The hourly demand for music downloads is given by:

Qd = 25 – 5P,

            where Qd is the number of songs demanded per hour, and is the price of downloading a song. 

Suppose that = $3. At this price, what does the price elasticity of demand equal? What is the total revenue?

At what price does ε = 1? What is the total revenue when the price is $2.50?

Explain whether your results to part a and part b are consistent with the Total Revenue Test.

 


1
Expert's answer
2022-03-09T10:59:59-0500

a)


elasticity of demand= slope of demand curve x P/Q

slope of demand curve=-5

"P=\$ 3"


"Q_d=25-5P"

"= 25-5(3)"


"=10"


elasticity of demand "=-5\times \frac{3}{10}"


"=-1.5"


b)

Total revenue is given by;


"P\times Q"


"P=\$3"


"Q_d=25-5P"


"= 25-5(3)"


"=10"


"Revenue=3\times 10"


"=\$30"


c)

The elasticity of demand= slope of demand curve x P/Q

PED=1


"1=5\times \frac {P}{Q}"


"\frac {P}{Q}=\frac {3}{10}"


"=0.3"


"\frac{P}{Q}=\frac{P}{25-5P}"


"0.3=\frac {P}{25-{5P}}"


"P=0.3(25-5P)"


"P=7.5-1.5P"


"P+1.5P=7.5"


"P=3"


Thus PED=1 when the price is $3


d)


Revenue= P X Q

When P=$2.5 Q will be:


"Q=25-5P"


"=25-5(2.5)"

"=12.5"


"2.5\times 12.5 =\$ 31.25"


e)

They are consistent with the total revenue test since a fall in price is leading to a rise in total revenue in the presence of elastic demand.


when the price falls from $3 to $2.5 the total revenue rises from $30 to $31.25. This is in line with the revenue test.


When a decrease in price causes an increase in total revenue, demand can be said to have been elastic since an increase in price has a greater impact on the quantity demanded.